# Review of deep learning models for crypto price prediction: implementation and evaluation

### v1

@@ Abstract
There has been much interest in accurate cryptocurrency price forecast models by investors and researchers. Deep Learning models are prominent machine learning techniques that have transformed various fields and have shown potential for finance and economics. Although various deep learning models have been explored for cryptocurrency price forecasting, it is not clear which models are suitable due to high market volatility. In this study, we review the literature about deep learning for cryptocurrency price forecasting and evaluate novel deep learning models for cryptocurrency stock price prediction. Our deep learning models include variants of long short-term memory (LSTM) recurrent neural networks, variants of convolutional neural networks (CNNs), and the Transformer model. We evaluate univariate and multivariate approaches for multi-step ahead predicting of cryptocurrencies close-price. Our results show that the univariate LSTM model variants perform best for cryptocurrency predictions. We also carry out volatility analysis on the four cryptocurrencies which reveals significant fluctuations in their prices throughout the COVID-19 pandemic. Additionally, we investigate the prediction accuracy of two scenarios identified by different training sets for the models. First, we use the pre-COVID-19 datasets to model cryptocurrency close-price forecasting during the early period of COVID-19. Secondly, we utilise data from the COVID-19 period to predict prices for 2023 to 2024.
@@ 3.5 Framework
In Step 4, we provide data analysis by first implementing a volatility analysis of the close price of selected cryptocurrencies, which can reveal fluctuations and patterns throughout the selected period. Historically, cryptocurrencies featured a wide price range, with high major fluctuation across the COVID-19 period [129, 130]. We use the volatility analysis to review the fluctuations, since they can lead to instability during the model training process, causing slower convergence and poor generalisation ability on the test dataset. In Step 4, taking into account our multivariate models, we provide feature correlation analysis to find out how the different features affect each other.
@@
In Step 6, we use Dataset 2 for Experiment 2 to predict the close price during COVID-19 using training data that features COVID-19 effect on the cryptocurrencies. We do this to determine whether the prediction accuracy has been improved and hence compare the results with Experiment 1. We also incorporate the shuffle data splitting strategy to enhance the efficacy of model performance. The initial 70% of the data is randomly rearranged using the shuffle.
@@ 4.2 Results: pre-COVID-19
Finally, we present the results for the Litcoin for both Univariate and Multivariate strategies. Figures 20 21 show all the results of the Univariate models, where we find that the Conv-LSTM and BD-LSTM show the best performance, whereas LSTM, ED-LSTM and Conv-LSTM provide the best performances for the Multivariate strategies. On the contrary, the CNN provides the worst performance for the Multivariate strategy, much higher in magnitude when compared to the rest of the models. We also notice that the Multivariate strategy provides better stability as the prediction horizon increases when compared to the Univariate strategy, and the Univariate strategy provides much better accuracy of the best models when compared to the Multivariate strategy. We summarise the results further in Table 10 which features the model prediction accuracy of the test dataset. We report the RMSE mean and 95% confidence interval for the four cryptocurrencies, and the best models for the different steps are highlighted in bold. The test mean provides the average of the five steps. It is clear that the Univariate models are better than the Multivariate models; however, we find that the accuracy of both strategies is close (test mean) for Bitcoin and Dogecoin.
@@
Table 10: Prediction accuracy on the test dataset reporting RMSE mean and 95% confidence interval ($\pm$) for the four cryptocurrencies. The best models for the different steps are highlighted in bold, and the test mean provides the average of the five steps.
@@ 4.3 Results: Data featuring COVID-19
The previous section presents results given by the respective models using data before COVID-19. We found that the univariate strategy was better than the Multivariate strategy (Table 10), therefore we only used the Univariate strategy for Experiment 2 (during COVID-19) and presented the results in Table 11. We find that comparing the results of Experiment 1, the prediction accuracy of Bitcoin, Ethereum, and Dogecoin has improved to a certain extent. The prediction accuracy that reveals the greatest improvement is forecasting the close price of Dogecoin. After training using data from the COVID-19 period, the prediction performance of Dogecoin is almost close to Bitcoin and Ethereum. Nevertheless, the forecast precision for Litecoin decreased. Additionally, we find that the robustness of the model improved after training with data from the COVID-19 period. The models keep that the confidence intervals for all predicted horizons of Bitcoin and Ethereum are controlled within $\pm$0.0007. In the case of Dogecoin and Litecoin, the robustness of the models in 1-step ahead prediction has generally improved.
Data Model Step 1 Step 2 Step 3 Step 4 Step 5 Test Mean
BTC BD-LSTM 0.0194$\pm$0.0002 0.0258$\pm$0.0003 0.0311$\pm$0.0004 0.0367$\pm$0.0003 0.0414$\pm$0.0004 0.0309$\pm$0.0003
@@
ETH BD-LSTM 0.0230$\pm$0.0002 0.0304$\pm$0.0004 0.0361$\pm$0.0005 0.0426$\pm$0.0004 0.0484$\pm$0.0006 0.0361$\pm$0.0004
@@
DOGE BD-LSTM 0.0291$\pm$0.0001 0.0618$\pm$0.0041 0.0673$\pm$0.0046 0.0742$\pm$0.0044 0.0795$\pm$0.0041 0.0624$\pm$0.0035
@@
LTC BD-LSTM 0.0577$\pm$0.0007 0.0797$\pm$0.0022 0.0968$\pm$0.0018 0.1096$\pm$0.0015 0.1205$\pm$0.0016 0.0929$\pm$0.0016
@@
Data Model LSTM ED-LSTM BD-LSTM CNN Conv-LSTM Transformer
@@
Multivariate 3 1 2 5 6 4
ETH Univariate 1 3 2 5 6 4
Multivariate 5 4 3 1 2 6
DOGE Univariate 4 3 1 6 2 5
Multivariate 2 3 1 6 4 5
LTC Univariate 4 3 2 5 1 6
Multivariate 1 3 5 6 2 4
Mean Rank 3 2.875 2.125 4.875 3.125 5
Table 12: Performance (rank) of different models for predicting cryptocurrency for Dataset 1.
Data LSTM ED-LSTM BD-LSTM CNN Conv-LSTM Transformer
BTC 4 3 1 5 2 6
@@
Table 13: Performance (rank) of different models for predicting cryptocurrency for Dataset 2.

### v2

@@ Abstract
There has been much interest in accurate cryptocurrency price forecast models by investors and researchers. Deep Learning models are prominent machine learning techniques that have transformed various fields and have shown potential for finance and economics. Although various deep learning models have been explored for cryptocurrency price forecasting, it is not clear which models are suitable due to high market volatility. In this study, we review the literature about deep learning for cryptocurrency price forecasting and evaluate novel deep learning models for cryptocurrency stock price prediction. Our deep learning models include variants of long short-term memory (LSTM) recurrent neural networks, variants of convolutional neural networks (CNNs), and the Transformer model. We evaluate univariate and multivariate approaches for multi-step ahead predicting of cryptocurrencies close-price. We also carry out volatility analysis on the four cryptocurrencies which reveals significant fluctuations in their prices throughout the COVID-19 pandemic. Additionally, we investigate the prediction accuracy of two scenarios identified by different training sets for the models. First, we use the pre-COVID-19 datasets to model cryptocurrency close-price forecasting during the early period of COVID-19. Secondly, we utilise data from the COVID-19 period to predict prices for 2023 to 2024. Our results show that the convolutional LSTM with a multivariate approach performs best. Our results also indicate that the multivariate deep learning models exhibit better performance in forecasting four different cryptocurrencies compared to the univariate models. Furthermore, the multivariate models show greater robustness in their prediction outcomes through numerous prediction horizons.
@@ 3.5 Framework
In Step 4, we provide data analysis by first implementing a volatility analysis of the close price of selected cryptocurrencies, which can reveal fluctuations and patterns throughout the selected period. Historically, cryptocurrencies featured a wide price range, with high major fluctuation across the COVID-19 period [129, 130]. We use the volatility analysis to review the fluctuations since they can lead to instability during the model training process, causing slower convergence and poor generalisation ability on the test dataset. In Step 4, taking into account our multivariate models, we provide feature correlation analysis to find out how the different features affect each other.
@@
In Step 6, we use Dataset 2 for Experiment 2 to predict the close price during COVID-19 using training data that features COVID-19 effect on the cryptocurrencies. We do this to determine whether the prediction accuracy has been improved and hence compare the results with Experiment 1. We also incorporate the shuffle data splitting strategy to enhance the efficacy of model performance. The initial 70% of the data is randomly rearranged using the shuffle strategy.
@@ 4.2 Results: pre-COVID-19
Finally, we present the results for the Litcoin for both Univariate and Multivariate strategies. Figures 20 and 21 show all the results of the Univariate models, where we find that the Conv-LSTM and BD-LSTM show the best performance, whereas LSTM, ED-LSTM and Conv-LSTM provide the best performances for the Multivariate strategies. On the contrary, the CNN provides the worst performance for the Multivariate strategy, much higher in magnitude when compared to the rest of the models. We also notice that the Multivariate strategy provides better stability as the prediction horizon increases, when compared to the Univariate strategy. The Univariate strategy provides much better accuracy of the best models when compared to the Multivariate strategy.
@@
Multivariate* CNN 0.0202$\pm$0.0007 0.0237$\pm$0.0008 0.0259$\pm$0.0007 0.0286$\pm$0.0008 0.0309$\pm$0.0008 0.0259$\pm$0.0008
LSTM 0.0178$\pm$0.0018 0.0213$\pm$0.0017 0.0241$\pm$0.0012 0.0269$\pm$0.0011 0.0294$\pm$0.0014 0.0237$\pm$0.0014
ED-LSTM 0.0171$\pm$0.0012 0.0214$\pm$0.0015 0.0240$\pm$0.0014 0.0269$\pm$0.0012 0.0294$\pm$0.0011 0.02378$\pm$0.0013
BD-LSTM 0.0175$\pm$0.0010 0.0210$\pm$0.0010 0.0239$\pm$0.0009 0.0263$\pm$0.0007 0.0291$\pm$0.0010 0.0236$\pm$0.0009
Conv-LSTM 0.0147$\pm$0.0004 0.0176$\pm$0.0004 0.0205$\pm$0.0004 0.0224$\pm$0.0006 0.0250$\pm$0.0006 0.0200$\pm$0.0005
Transformer 0.0186$\pm$0.0008 0.0210$\pm$0.0008 0.0231$\pm$0.0008 0.0257$\pm$0.0008 0.0279$\pm$0.0008 0.0233$\pm$0.0008
@@
Multivariate* CNN 0.0202$\pm$0.0005 0.0228$\pm$0.0007 0.0243$\pm$0.0006 0.0266$\pm$0.0008 0.0282$\pm$0.0010 0.0244$\pm$0.0007
LSTM 0.0168$\pm$0.0009 0.0217$\pm$0.0014 0.0241$\pm$0.0007 0.0274$\pm$0.0010 0.0311$\pm$0.0014 0.0242$\pm$0.0011
ED-LSTM 0.0179$\pm$0.0010 0.0223$\pm$0.0013 0.0257$\pm$0.0011 0.0278$\pm$0.0013 0.0308$\pm$0.0013 0.0249$\pm$0.0012
BD-LSTM 0.0171$\pm$0.0008 0.0210$\pm$0.00006 0.0246$\pm$0.0009 0.0278$\pm$0.0010 0.0300$\pm$0.0011 0.0241$\pm$0.0009
Conv-LSTM 0.0158$\pm$0.0006 0.0193$\pm$0.0006 0.0212$\pm$0.0008 0.0231$\pm$0.0007 0.0255$\pm$0.0009 0.0210$\pm$0.0007
Transformer 0.0186$\pm$0.0009 0.0201$\pm$0.0009 0.0220$\pm$0.0008 0.0239$\pm$0.0010 0.0262$\pm$0.0009 0.0222$\pm$0.0009
Table 10: Prediction accuracy on the test dataset (RMSE mean and 95% confidence interval ($\pm$)) for Bitcoin and Ethereum. The best models for the different steps are highlighted in bold, and the test mean provides the average of the five steps. Note that the Multivariate strategy only features Gold price and features (high, opening, low) to predict the close price of the given cryptocurrency (e.g. Bitcoin). In the Multivariate* strategy, there is an addition of the close price of the most correlated cryptocurrency (e.g. Bitcoin uses Ethereum as shown in Figure 13)
Data Strategy Model Step 1 Step 2 Step 3 Step 4 Step 5 Test Mean
@@
Multivariate* CNN 0.0220$\pm$0.0017 0.0257$\pm$0.0017 0.0301$\pm$0.0019 0.0326$\pm$0.0022 0.0330$\pm$0.0022 0.0287$\pm$0.0019
LSTM 0.0194$\pm$0.0013 0.0238$\pm$0.0015 0.0260$\pm$0.0011 0.0307$\pm$0.0017 0.0323$\pm$0.0008 0.0264$\pm$0.0013
ED-LSTM 0.0196$\pm$0.0020 0.0248$\pm$0.0022 0.0281$\pm$0.0023 0.0317$\pm$0.0025 0.0348$\pm$0.0021 0.0278$\pm$0.0022
BD-LSTM 0.0183$\pm$0.0013 0.0234$\pm$0.0013 0.0272$\pm$0.0015 0.0317$\pm$0.0018 0.0350$\pm$0.0021 0.0271$\pm$0.0016
Conv-LSTM 0.0168$\pm$0.0014 0.0173$\pm$0.0012 0.0179$\pm$0.0015 0.0200$\pm$0.0018 0.0230$\pm$0.0018 0.0190$\pm$0.0015
Transformer 0.0208$\pm$0.0016 0.0212$\pm$0.0015 0.0237$\pm$0.0016 0.0271$\pm$0.0019 0.0304$\pm$0.0022 0.0246$\pm$0.0018
@@
Multivariate* CNN 0.0277$\pm$0.0011 0.0304$\pm$0.0012 0.0342$\pm$0.0012 0.0371$\pm$0.0011 0.0409$\pm$0.0011 0.0341$\pm$0.0011
LSTM 0.0233$\pm$0.0015 0.0295$\pm$0.0016 0.0340$\pm$0.0016 0.0380$\pm$0.0017 0.0416$\pm$0.0017 0.0333$\pm$0.0016
ED-LSTM 0.0230$\pm$0.0012 0.0300$\pm$0.0015 0.0350$\pm$0.0016 0.0395$\pm$0.0014 0.0433$\pm$0.0016 0.0342$\pm$0.0015
BD-LSTM 0.0240$\pm$0.0013 0.0306$\pm$0.0010 0.0355$\pm$0.0010 0.0395$\pm$0.0013 0.0432$\pm$0.0013 0.0346$\pm$0.0012
Conv-LSTM 0.0223$\pm$0.0008 0.0270$\pm$0.0009 0.0307$\pm$0.0012 0.0332$\pm$0.0013 0.0362$\pm$0.0011 0.0300$\pm$0.0011
Transformer 0.0253$\pm$0.0013 0.0276$\pm$0.0012 0.0303$\pm$0.0008 0.0336$\pm$0.0010 0.0374$\pm$0.0014 0.0308$\pm$0.0011
Table 11: Prediction accuracy for the test dataset (RMSE mean and 95% confidence interval ($\pm$)) for Dodgecoin and Litecoin. The best models for the different steps are highlighted in bold, and the test mean provides the average of the five steps. Note that the Multivariate strategy only features Gold price and features (high, opening, low) to predict the close price of the given cryptocurrency (e.g. Dogecoin). In the Multivariate* strategy, there is an addition of the close price of the most correlated cryptocurrency (e.g. Dodgecoin uses Ethereum as shown in Figure 13).
Furthermore, we present another approach (Multivariate*) which features the Gold price and a highly correlated cryptocurrency as shown in Figure 13, i.e. in the case of Bitcoin, the Multivariate* model features the Ethereum and Gold price along with Bitcoin high, low, opening and closing price. Figure 22 presents the results where we observe that the forecasting results improve significantly given the additional feature of a highly correlated cryptocurrency, when compared to previous Multivariate results for the respective cryptocurrencies (Figure 15, 17, 19, and 21). In Figure 22, the prediction accuracy of all the respective models exhibit a consistent decrease in accuracy (higher RMSE) with increasing prediction horizon. The prediction results for Bitcoin shows that Conv-LSTM is the best performing model in Figure 22a and CNN is the worst model. Figure 22b shows the results of each model predicting Ethereum where Conv-LSTM maintains the best prediction accuracy. It is worth to mention that the ranking of LSTM among the 6 models gradually deteriorates as the prediction horizon increases. The prediction results of the models for Dogecoin are shown in Figure 22c where Conv-LSTM has the best prediction performance in this case, followed by the Transformer model. Figure 22d presents the results of the Multivariate* for Litecoin where we find that Conv-LSTM shows the best prediction accuracy.
We summarise the results further in Tables 10 and 11 which feature the model prediction accuracy of the test dataset. We report the RMSE mean and 95% confidence interval for the four cryptocurrencies, and the best models for the different steps are highlighted in bold. The test mean provides the average of the five steps. It is clear that the Univariate models are better than the Multivariate models; however, we find that the accuracy of both strategies is close (test mean) for Bitcoin and Dogecoin. In general, Multivariate* strategy provides the best results for the four cryptocurrencies.
Table 13 provides a rank of the models where BD-LSTM provides the best accuracy rank followed by Conv-LSTM for the case of the Univariate approach. Furthermore, Conv-LSTM provides the best accuracy rank followed by the LSTM and ED-LSTM models for the Multivariate approach. Finally, Conv-LSTM provides the best accuracy rank followed by the Transformer model for the Multivariate* approach.
(a) Bitcoin
(b) Ethereum
(c) Dogecoin
(d) Litecoin
Figure 22: Performance evaluation of respective Multivariate* methods (RMSE mean with 95% CI for 30 experimental runs.) for Experiment 1: pre COVID-19.
@@ 4.3 Results: Data featuring COVID-19
The previous section presented results given by the respective models using data before COVID-19. We found that the Multivariate* strategy provided the best performance (Table 10 and 11) in comparison to Multivariate and Univariate strategies. Therefore, in Experiment 2, i.e. data featuring COVID-19, we only used the Univariate and Multivariate* strategies and presented the results in Table 12. We further present the results for predicting the cryptocurrency at each prediction horizon in Figure 23. We find that Conv-LSTM is the best model under the Multivariate* strategy. Figure 23a displays the performance of 6 deep learning models in forecasting Bitcoin. As the number of prediction days increases, the forecast accuracy gradually decreases (higher RMSE). The models predict that the result of Ethereum (in Figure 23b) is similar to that of Bitcoin. Figure 23c demonstrates a decline in the model’s robustness when making predictions for Dogecoin. The 95% confidence interval of the model is larger than the other three cryptocurrencies. In terms of prediction accuracy, Dogecoin is almost the same as other cryptocurrencies. We also present the forecast results of Litecoin in Figure 23d, although the outcomes remain consistent with Bitcoin, Ethereum and Dogecoin.
We find that comparing the results of Experiment 2 with Experiment 1 (pre-COVID-19 data), the prediction accuracy of Bitcoin, Ethereum, and Dogecoin has generally deteriorated to a certain extent. Let us examine the case of Conv-LSTM with Multivariate* which has provided the best results for Experiment 1 (Tables 10 and 11) and Experiment 2 (Table 12). Looking at the test mean, we get RMSE around 0.02 for Bitcoin and Ethereum in Experiment 1, but in Experiment 2, we get RMSE of around 0.03. Furthermore, in Experiment 1 Multivariate* Conv-LSTM, we find that Dogecoin and Litecoin obtained RMSE of 0.019 and 0.03, respectively; in Experiment 2, these deteriorated to 0.029 and 0.032, respectively. This could be due to the high vitality in the dataset featuring COVID-19 as the markets were adjusting due to lockdowns and economic crisis around the world.
Table 14 provides a rank of the models where BD-LSTM provides the best accuracy rank followed by ED-LSTM for the Univariate strategy. Furthermore, Conv-LSTM provides the best accuracy rank followed by the BD-LSTM models for the Multivariate* approach.
Data Strategy Model Step 1 Step 2 Step 3 Step 4 Step 5 Test Mean
BTC Univariate BD-LSTM 0.0194$\pm$0.0002 0.0258$\pm$0.0003 0.0311$\pm$0.0004 0.0367$\pm$0.0003 0.0414$\pm$0.0004 0.0309$\pm$0.0003
@@
Multivariate* BD-LSTM 0.0226$\pm$0.0009 0.0286$\pm$0.0008 0.0335$\pm$0.0007 0.0378$\pm$0.0009 0.0421$\pm$0.0008 0.0329$\pm$0.0008
ED-LSTM 0.0237$\pm$0.0014 0.0304$\pm$0.0017 0.0349$\pm$0.0016 0.0392$\pm$0.0014 0.0437$\pm$0.0015 0.0344$\pm$0.0015
LSTM 0.0235$\pm$0.0012 0.0295$\pm$0.0013 0.0339$\pm$0.0010 0.0380$\pm$0.0010 0.0426$\pm$0.0013 0.0335$\pm$0.0012
CNN 0.0289$\pm$0.0010 0.0336$\pm$0.0009 0.0375$\pm$0.0009 0.0411$\pm$0.0010 0.0446$\pm$0.0008 0.0371$\pm$0.0009
Conv-LSTM 0.0220$\pm$0.0007 0.0276$\pm$0.0006 0.0323$\pm$0.0006 0.0368$\pm$0.0007 0.0405$\pm$0.0006 0.0318$\pm$0.0006
Transformer 0.0265$\pm$0.0010 0.0311$\pm$0.0009 0.0350$\pm$0.0008 0.0387$\pm$0.0010 0.0423$\pm$0.0008 0.0347$\pm$0.0009
ETH Univariate BD-LSTM 0.0230$\pm$0.0002 0.0304$\pm$0.0004 0.0361$\pm$0.0005 0.0426$\pm$0.0004 0.0484$\pm$0.0006 0.0361$\pm$0.0004
@@
Multivariate* BD-LSTM 0.0245$\pm$0.0011 0.0307$\pm$0.0012 0.0354$\pm$0.0012 0.0403$\pm$0.0015 0.0442$\pm$0.0014 0.0350$\pm$0.0013
ED-LSTM 0.0245$\pm$0.0011 0.0314$\pm$0.0014 0.0358$\pm$0.0014 0.0404$\pm$0.0013 0.0443$\pm$0.0013 0.0353$\pm$0.0013
LSTM 0.0254$\pm$0.0015 0.0313$\pm$0.0017 0.0362$\pm$0.0014 0.0404$\pm$0.0012 0.0445$\pm$0.0014 0.0356$\pm$0.0014
CNN 0.0296$\pm$0.0007 0.0337$\pm$0.0008 0.0386$\pm$0.0009 0.0413$\pm$0.0010 0.0456$\pm$0.0010 0.0378$\pm$0.0009
Conv-LSTM 0.0228$\pm$0.0008 0.0284$\pm$0.0007 0.0324$\pm$0.0008 0.0365$\pm$0.0010 0.0398$\pm$0.0009 0.0320$\pm$0.0008
Transformer 0.0287$\pm$0.0014 0.0328$\pm$0.0015 0.0366$\pm$0.0.014 0.0399$\pm$0.0016 0.0438$\pm$0.0015 0.0364$\pm$0.0015
DOGE Univariate BD-LSTM 0.0291$\pm$0.0001 0.0618$\pm$0.0041 0.0673$\pm$0.0046 0.0742$\pm$0.0044 0.0795$\pm$0.0041 0.0624$\pm$0.0035
@@
Multivariate* BD-LSTM 0.0238$\pm$0.0015 0.0293$\pm$0.00021 0.0349$\pm$0.0021 0.0404$\pm$0.0022 0.0446$\pm$0.0024 0.0346$\pm$0.0021
ED-LSTM 0.0254$\pm$0.0018 0.0313$\pm$0.0024 0.0374$\pm$0.0020 0.0428$\pm$0.0025 0.0464$\pm$0.0022 0.0367$\pm$0.0022
LSTM 0.0240$\pm$0.0020 0.0301$\pm$0.0018 0.0345$\pm$0.0016 0.0406$\pm$0.0017 0.0449$\pm$0.0019 0.0348$\pm$0.0018
CNN 0.0275$\pm$0.0017 0.0327$\pm$0.0022 0.0377$\pm$0.0017 0.0413$\pm$0.0020 0.0449$\pm$0.0030 0.0368$\pm$0.0021
Conv-LSTM 0.0236$\pm$0.0020 0.0266$\pm$0.0020 0.0286$\pm$0.0019 0.0327$\pm$0.0023 0.0375$\pm$0.0026 0.0298$\pm$0.0022
Transformer 0.0323$\pm$0.0023 0.0340$\pm$0.0023 0.0375$\pm$0.0022 0.0411$\pm$0.0022 0.0460$\pm$0.0023 0.0382$\pm$0.0023
LTC Univariate BD-LSTM 0.0577$\pm$0.0007 0.0797$\pm$0.0022 0.0968$\pm$0.0018 0.1096$\pm$0.0015 0.1205$\pm$0.0016 0.0929$\pm$0.0016
@@
Multivariate* BD-LSTM 0.0247$\pm$0.0012 0.0306$\pm$0.0010 0.0356$\pm$0.0013 0.0401$\pm$0.0013 0.0453$\pm$0.0014 0.0352$\pm$0.0012
ED-LSTM 0.0249$\pm$0.0015 0.0320$\pm$0.0020 0.0369$\pm$0.0019 0.0413$\pm$0.0020 0.0457$\pm$0.0019 0.0362$\pm$0.0019
LSTM 0.0254$\pm$0.0014 0.0310$\pm$0.0012 0.0361$\pm$0.0014 0.0399$\pm$0.0016 0.0454$\pm$0.0018 0.0357$\pm$0.0015
CNN 0.0304$\pm$0.0010 0.0348$\pm$0.0011 0.0396$\pm$0.0012 0.0434$\pm$0.0011 0.0466$\pm$0.0012 0.0390$\pm$0.0011
Conv-LSTM 0.0239$\pm$0.0009 0.0290$\pm$0.0010 0.0325$\pm$0.0008 0.0360$\pm$0.0010 0.0403$\pm$0.0012 0.0323$\pm$0.0010
Transformer 0.0295$\pm$0.0015 0.0331$\pm$0.0016 0.0367$\pm$0.0018 0.0394$\pm$0.0019 0.0434$\pm$0.0017 0.0364$\pm$0.0017
@@
(a) Bitcoin
(b) Ethereum
(c) Dogecoin
(d) Litecoin
Figure 23: Performance evaluation of respective Multivariate* methods (RMSE mean with 95% CI for 30 experimental runs.) for Experiment 2: data featuring COVID-19.
Data Strategy LSTM ED-LSTM BD-LSTM CNN Conv-LSTM Transformer
@@
ETH 1 3 2 5 6 4
DOGE 4 3 1 6 2 5
LTC 4 3 2 5 1 6
Mean Rank 3.25 3 1.5 5.25 2.75 5.25
BTC Multivariate 3 1 2 5 6 4
ETH 5 4 3 1 2 6
DOGE 2 3 1 6 4 5
LTC 1 3 5 6 2 4
Mean Rank 2.75 2.75 2.75 4.5 3.5 4.75
BTC Multivariate* 4 5 3 6 1 2
ETH 4 6 3 5 1 2
DOGE 3 5 4 6 1 2
LTC 3 5 6 4 1 2
Mean Rank 3.5 5.25 4 5.25 1 2
Table 13: Performance (rank) of different models for predicting cryptocurrency for Experiment 1.
Data Strategy LSTM ED-LSTM BD-LSTM CNN Conv-LSTM Transformer
BTC Univariate 4 3 1 5 2 6
@@
BTC Multivariate* 3 4 2 6 1 5
ETH 4 3 2 6 1 5
DOGE 3 4 2 5 1 6
LTC 3 4 2 6 1 5
Mean Rank 3.25 3.75 2 5.75 1 5.25
Table 14: Performance (rank) of different models for predicting cryptocurrency for Experiment 2.